OpenStudy (anonymous):

for rectangle ABCD, if AB=14 and EB=10, then what is the measure of AC?

OpenStudy (raden):

do you have a diagram, just want know where the position of point E ?

OpenStudy (anonymous):

look at number 14 on the pic i just sent

OpenStudy (raden):

nah, that make sense now. obvious, the diagonals of a rectangle always be same, so AC = BD. while BD = 2 * EB = 2 * 5 = 10. therefore AC = 10